# Conceptualization, Formal analysis, Investigation, Methodology, Software, Data curation, Validation, Visualization, Writing –original draft, Writing – review & editing

### v3

@@ 3.6 Attention-based Refinement Process
where $\mathbf{H}$ is the calibration matrix from earlier stages, and $\mathbf{H}^{\Delta}$ compensates for non-planar distortions, lens imperfections, and other real-world discrepancies.
  1. 1.
    Vision Transformer for Global Distortion Features: Unlike purely convolutional networks, a Vision Transformer (ViT) employs self-attention to capture global cues from the input image. Let $\mathbf{I}$ denote the image, partitioned into patches $\{\mathbf{p}_{1},\dots,\mathbf{p}_{m}\}$. The ViT encodes each patch as a token $\mathbf{t}_{i}=f_{\mathrm{ViT}}(\mathbf{p}_{i})$, and uses multi-head attention to capture long-range dependencies:
    $\mathrm{Attn}\!\bigl{(}\mathbf{Q},\mathbf{K},\mathbf{V}\bigr{)}\;=\;\mathrm{Softmax}\!\Bigl{(}\frac{\mathbf{Q}\,\mathbf{K}^{\top}}{\sqrt{d}}\Bigr{)}\,\mathbf{V},$
    where $d$ is the token dimension. By aggregating information across patches, the ViT reveals global distortion features that a simple planar model cannot accommodate. This capability helps mitigate inaccuracies stemming from homography’s planar assumptions or the absence of camera-intrinsic rectification.
  2. 2.
    Cross-Attention for LiDAR–Camera Correspondences: To establish more accurate correspondences between LiDAR and camera data, we integrate a cross-attention mechanism that takes:
    • Queries $\mathbf{Q_{c}}$ from image points, $\{(u_{i},v_{i})\}_{i=1}^{N}$.
    • Keys $\mathbf{K_{c}}$ from projected LiDAR points, $\{\hat{u}_{j},\hat{v}_{j}\}_{j=1}^{M}$, derived by transforming each LiDAR 2D coordinate $(x_{j},y_{j})$ via $\mathbf{H}$.
    • Values $\mathbf{V_{c}}$ from the original LiDAR points, $\{(x_{j},y_{j},z_{j})\}_{j=1}^{M}$, retaining full 3D spatial information.
    By computing weighted interactions $\mathbf{A}_{\mathrm{cross}}$ between these queries and keys,
@@
    the model effectively links each image point with its corresponding 3D LiDAR counterpart. This cross-modal attention captures more accurate 2D–3D relationships and refines the camera–LiDAR alignment beyond what a purely homography-based approach can achieve.
  3. 3.
    Learning the Correction Matrix: Let $\mathbf{z}_{\mathrm{ViT}}$ be the output tokens from the Vision Transformer and $\mathbf{A}_{\mathrm{cross}}$ be the cross-attention output. We concatenate these to form a feature vector
@@
    which is passed through additional layers $g(\cdot)$ that regress a 9-dimensional vector $\boldsymbol{\theta}$, representing the correction matrix:
    $\boldsymbol{\theta}=g\!\bigl{(}\mathbf{f}\bigr{)},\quad\mathbf{H}^{\Delta}=\mathrm{Reshape}\!\bigl{(}\boldsymbol{\theta}\bigr{)}\;\in\;\mathbb{R}^{3\times 3}.$
    The final refined homography is
    $\mathbf{H}^{*}=\mathbf{H}\,\times\,\mathbf{H}^{\Delta}.$
    During training, a self-supervised objective seeks to minimize the reprojection error between the $\mathbf{H}^{*}$-projected LiDAR points and their corresponding image coordinates, thus driving $\mathbf{H}^{\Delta}$ to compensate for any non-planar misalignments left by the base homography.
@@
4 Experimental Results and Analysis
4.1 Sensor Setup and Data Collection
@@
Dataset 1: The first dataset was collected at the intersection of M.L.K. Boulevard and Central Avenue in the Chattanooga Smart Corridor, where a comprehensive two-hour synchronized dataset was gathered using multiple sensor types. The sensor setup consisted of 32-channel LiDAR systems mounted on utility poles at the intersection corners, as illustrated in Fig. 7. These LiDAR units operate with a detection range of 0.05 to 120 meters and are complemented by integrated video cameras. The dual-sensor approach combines precise LiDAR spatial measurements with video footage, providing both quantitative data and visual context for validating observations.
@@
Dataset 2. The second dataset was collected at an urban intersection in downtown Chattanooga, specifically at Georgia Avenue and M.L.K. Boulevard. This site was chosen for its high traffic volume and diverse mix of road users. Similar to Dataset 1, a LiDAR–camera system was employed to collect the data; LiDAR scans and camera images were synchronized via ROS and stored in ROSbag files with precise timestamps. This setup ensures accurate alignment of multi-modal data, thereby facilitating the investigation of cross-sensor calibration.
4.2 Deep Learning Model Training
4.2.1 Data Annotation and Dataset Generation
We developed a multi-sensor annotation toolkit designed for efficiently labeling common objects captured by both camera images and point cloud data, with applications in autonomous driving and intelligent transportation research. This toolkit integrates both automatic and manual annotation methods, thereby optimizing the labeling process while ensuring high-quality datasets.
For bounding box generation, the toolkit employs two primary approaches. In camera image annotation, it first utilizes a YOLO-based object detection algorithm to automatically generate bounding boxes and then allows users to manually draw boxes for precise labeling. For LiDAR point cloud data, the toolkit automatically extracts the background through preprocessing algorithms and uses DBSCAN clustering to detect objects and generate initial bounding boxes, which can subsequently be refined manually if necessary.
Furthermore, once detections from both camera and LiDAR are available, the toolkit offers a dual-view visualization interface that enables users to match identical objects across the two sensor modalities. Using this system, we annotated the two datasets described above: Dataset 1 consists of 1200 frames, of which 800 frames were annotated, resulting in a total of 5,815 labeled identical objects; Dataset 2 consists of 600 frames, with 200 frames annotated and a total of 619 labeled identical objects.
4.2.2 Training Details
@@
The Common Feature Discriminator and the Attention-based Model were both trained from scratch on the UArizona High-Performance Computing Platform, utilizing a computational setup that included a single Nvidia 32GB V100S GPU and an AMD Zen2 processor with 5 cores and 30GB of RAM. The training was implemented using PyTorch 2.0 as the deep learning framework. The optimization process was carried out using the Adam optimizer, configured with a momentum of 0.937 and a weight decay of $5\times 10^{-4}$. A cosine learning rate schedule was employed to dynamically adjust the learning rate, beginning with an initial learning rate of 0.001, warming up with a ratio of 0.05, and gradually decaying to a minimum learning rate of 0.00001. To further enhance training stability, an exponential moving average (EMA) strategy with a decay rate of 0.9999 was adopted.
Both datasets were divided into training, validation, and test sets, with 90% allocated to the training-validation split and 10% reserved for testing. Within the training-validation split, 90% was designated for training and 10% for validation. Both the training and validation sets were shuffled before each epoch to ensure robust learning. No data augmentation techniques were used, except for resizing to fit the model requirements. For the Common Feature Discriminator, training was conducted over 300 epochs with a batch size of 4. For the Attention-based Model, the training spanned 800 epochs with a batch size of 8, and the token length was set to 256.
4.3 Results and Discussion
@@
       1) Common Feature Discriminator Performance:
Table 1 summarizes the Common Feature Discriminator’s performance on both datasets, revealing consistently strong binary classification accuracies (98.00% and 92.50% for Datasets 1 and 2, respectively). These high scores indicate that the discriminator is highly effective at distinguishing whether pairs of LiDAR and camera detections correspond to the same physical object. Meanwhile, the slightly lower image and LiDAR classification accuracies reflect the inherent variability in each modality’s appearance and point cloud density, as well as the increased complexity of Dataset 2’s urban traffic scenes. Overall, the results confirm that the discriminator robustly balances object-level matching (binary classification) with semantic recognition (modality-specific class labels). Figure 6 further illustrates the model’s qualitative behavior: two distinct objects (“differ”) are correctly identified as different, while two identical objects from different sensor views are consistently classified as “same.” This underlines the model’s robustness when handling variations in object types and poses. Notably, though occasional misclassifications occur—such as trucks being predicted as cars—these errors are relatively rare and do not significantly affect the system’s ability to produce reliable point correspondences.
Table 2: Performance Comparison of Coarse Calibration and Other Methods
@@
       2) Coarse Calibration Accuracy:
Once the Common Feature Discriminator identifies matching objects across LiDAR and camera modalities and the corresponding point pairs are derived (as shown in Fig. 8(a)), Homography calibration is employed to obtain a coarse calibration matrix. To comprehensively evaluate the accuracy of this coarse solution, we compare it with several existing calibration methods: Manual calibration, LCCNet (Lv et al., 2021), CalibDepth (Zhu et al., 2023), and CalibNet (Iyer et al., 2018). It is worth noting that for Manual calibration, we selected 34 representative point pairs uniformly distributed across the sensors’ fields of view through manual object matching. Due to the time-intensive nature of this process, we did not exhaustively select all possible point pairs. Thus, the manual calibration results presented here represent a typical calibration effort within a reasonable timeframe, rather than a full-effort exhaustive manual optimization. Table 2 presents the results in terms of the reprojection error metrics $\mathcal{E}_{\text{AED}}$ and $\mathcal{E}_{\text{RMSE}}$ (defined in Section 3.1). From Table 2, our coarse calibration demonstrates comparable or, in some cases, superior performance compared to other methods. Specifically, the automated coarse calibration outperforms Manual calibration on Dataset 2, although it exhibits a slightly higher reprojection error than Manual calibration on Dataset 1 (as visualized in Fig. 8(b)). Nevertheless, the significant advantages of the automated approach in real-time operation and reduced human intervention render this trade-off both acceptable and practical. Furthermore, while the coarse method occasionally exhibits slightly higher errors than certain deep learning–based solutions (e.g., LCCNet), it consistently surpasses others (e.g., CalibDepth and CalibNet), underscoring the effectiveness of the proposed strategy. Fig. 9 presents example calibration outcomes, with red dots (camera detections) and green dots (LiDAR detections) projected onto the image plane. Despite some minor misalignments—particularly in the upper portion of the scene (highlighted by the yellow ellipse)—the coarse calibration overall provides a notably tighter alignment between the two sensor views, potentially enabling precise LiDAR-camera fusion in real-world traffic scenarios.
Table 3: Reprojection Error Evolution in Iterative Refinement for Dataset 1
@@
4.3.1 Fine Calibration with Iterative Refinement
@@
Table 4: Reprojection Error Evolution in Iterative Refinement for Dataset 2
@@
Tables 3 and 4 detail the reprojection error evolution (using the $\mathcal{E}_{\text{AED}}$ metric) at different frame intervals (with an interval of 100 frames in our implementation) for Datasets 1 and 2. In each interval, the algorithm determines whether the newly computed homography matrix (New) provides a tighter alignment than the previously best-known matrix (Best); if so, it updates the calibration accordingly. Fig. 10(a) and 10(b) visualize these updates, where the blue line denotes the error obtained from the newly recalibrated matrix in each iteration, and the orange line tracks the evolving best-known solution. Not every recalibration step yields an improvement—reflecting the inherent noise and variability of real-world data—but key frame intervals (e.g., 0–100 for Dataset 1 and 0–1000 for Dataset 2) demonstrate significant error reductions, confirming that the iterative approach converges toward a more accurate solution over time. These updates demonstrate the iterative optimization process’s ability to adaptively refine the calibration as additional data and correspondences become available, ultimately enabling the iterative refinement to achieve significantly higher accuracy compared to the initial coarse calibration (as shown in Fig. 12(b)).
Table 5: Performance Comparison of Iterative Refined Calibration and Other Methods
@@
Table 5 compares the final refined calibration performance with the aforementioned methods. Notably, the iterative refinement outperforms manual calibration by a sizeable margin in both datasets, reducing $\mathcal{E}_{\text{AED}}$ from 131.04 to 95.89 in Dataset 1 and from 40.79 to 28.35 in Dataset 2. It also consistently surpasses CalibDepth and CalibNet, while maintaining a competitive edge against LCCNet. These results demonstrate the effectiveness of iteratively incorporating new point correspondences in mitigating decalibrations and refining the sensor alignment. In practice, the procedure offers a compelling balance between accuracy, adaptability, and reduced reliance on strictly supervised or fully manual calibration protocols—making it especially valuable in long-term deployment scenarios.
Figure 13: Example Images demonstrating the Performance Improvement of Attention-based Refinement compared to Iterative Refinement.
@@
4.3.2 Fine Calibration with Attention-based Refinement
Although the iterative refinement approach already demonstrates strong performance, it remains inherently limited by the planar assumptions of Homography. Our proposed attention-based refinement aims to mitigate errors caused by image distortions and non-planar surfaces. As shown in Table 6, calibration after applying attention-based refinement (i.e., CalibRefine) achieves lower reprojection errors than other methods on both datasets, surpassing the iterative refinement (Table 5) in most metrics. Fig. 12(c) offers a more granular view of these improvements by comparing the $\mathcal{E}_{\text{AED}}$ distributions of iterative refinement and attention-based refinement. While the latter still exhibits some overlap with the former, its overall distribution skews toward smaller errors, indicating a more consistently accurate alignment between LiDAR and camera data. Fig. 13 visually illustrates such performance gains of attention-based refinement over iterative refinement.
Table 6: Performance Comparison of CalibRefine and Other Methods
@@ 4.3.2 Fine Calibration with Attention-based Refinement
Overall, our proposed CalibRefine framework consolidates three core components—Coarse Calibration, Iterative Refinement, and Attention-Based Refinement—into a unified solution. As illustrated in Fig. 12, each stage progressively refines the LiDAR–camera alignment, mitigating errors introduced by imperfect correspondence matching (coarse stage), limited point redundancy (iterative stage), or planar homography assumptions (attention-based stage). Table 6 further demonstrates that CalibRefine surpasses existing state-of-the-art methods in terms of quantitative reprojection accuracy. Beyond numerical metrics, Fig. 15 and 16 offer visual validation on Datasets 1 and 2, respectively, revealing how CalibRefine more reliably overlays LiDAR points with their corresponding image objects—particularly at scene edges and larger distances. In addition, Fig. 17 examines the distribution of calibration errors ($\mathcal{E}_{\text{AED}}$ and $\mathcal{E}_{\text{RMSE}}$) across competing approaches. Not only does CalibRefine exhibit a lower median error, but the overall spread of high-error outliers is also reduced, indicating its consistent performance. These findings underscore the robustness and adaptability of CalibRefine in real-world traffic environments.
5 Conclusion
In this paper, we presented CalibRefine, an end-to-end, fully automatic, targetless, and online LiDAR–camera calibration framework that integrates three core steps—coarse calibration, iterative refinement, and attention-based refinement—into a unified pipeline. By combining robust object detection with a Common Feature Discriminator, our method circumvents the need for manually placed fiducials or human-labeled sensor parameters. The coarse calibration phase provides a strong initial alignment, which the iterative refinement then continuously improves by leveraging newly acquired point correspondences across frames. Finally, the attention-based stage applies a Vision Transformer and cross-attention to handle non-planar distortions and subtle mismatches beyond the scope of homography.
Experiments on real-world urban datasets confirm that CalibRefine achieves accurate sensor alignment comparable to, and often better than, existing methods. Moving forward, the approach could benefit from exploring more advanced deep learning architectures or sophisticated mapping mechanisms, as well as extending the attention mechanism to incorporate scene geometry. Such enhancements could enable even more precise and high-fidelity calibration, particularly in large-scale deployment scenarios.

### v4

@@ III-F Attention-based Refinement Process
where $\mathbf{H}$ is the initial calibration matrix and $\mathbf{H}^{\Delta}$ compensates for non-planar distortions, lens imperfections, and other real-world discrepancies.
Our approach leverages a Vision Transformer (ViT) to capture global distortion features. Given an image $\mathbf{I}$ partitioned into patches $\{\mathbf{p}_{1},\dots,\mathbf{p}_{m}\}$, each patch is encoded into a token $\mathbf{t}_{i}=f_{\mathrm{ViT}}(\mathbf{p}_{i})$. The ViT applies multi-head self-attention,
$\mathrm{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})\;=\;\mathrm{Softmax}\!\Bigl{(}\frac{\mathbf{Q}\,\mathbf{K}^{\top}}{\sqrt{d}}\Bigr{)}\mathbf{V},$
with token dimension $d$, thereby aggregating global information to reveal distortion patterns that simple planar models cannot capture.
Simultaneously, a cross-attention mechanism establishes precise correspondences between LiDAR and camera data. Image points $\{(u_{i},v_{i})\}_{i=1}^{N}$ generate queries $\mathbf{Q_{c}}$, while LiDAR points projected by $\mathbf{H}$ yield keys $\mathbf{K_{c}}$ and the original 3D coordinates $\{(x_{j},y_{j},z_{j})\}_{j=1}^{M}$ serve as values $\mathbf{V_{c}}$. The cross-attention output is computed as
@@
effectively linking each image point with its corresponding LiDAR feature.
The global features from the ViT and the detailed correspondence information from cross-attention are concatenated to form a unified feature vector,
@@
which is then processed by additional layers $g(\cdot)$ to regress a 9-dimensional vector $\bm{\theta}$. This vector is reshaped to obtain the correction matrix,
$\bm{\theta}=g\!\bigl{(}\mathbf{f}\bigr{)},\ \ \mathbf{H}^{\Delta}\;=\;\mathrm{Reshape}\!\bigl{(}\bm{\theta}\bigr{)}\in\mathbb{R}^{3\times 3},$
resulting in the final refined homography $\mathbf{H}^{*}=\mathbf{H}\times\mathbf{H}^{\Delta}$. During training, a self-supervised loss minimizes the reprojection error between $\mathbf{H}^{*}$-projected LiDAR points and their corresponding image coordinates, guiding $\mathbf{H}^{\Delta}$ to correct any residual misalignments.
@@
IV Experimental Results and Analysis
IV-A Sensor Setup and Data Collection
@@
Dataset 1 was collected at the intersection of M.L.K. Boulevard and Central Avenue in the Chattanooga Smart Corridor, where a two-hour synchronized dataset was gathered using multiple sensor types. A 32-channel LiDAR system was mounted on utility poles at the intersection corners (Fig. 13), operating with a detection range of 0.05–120 m and complemented by integrated video cameras.
@@
Dataset 2 was collected at another urban intersection in downtown Chattanooga (Georgia Avenue and M.L.K. Boulevard), also employing a LiDAR–camera system. LiDAR scans and camera images were synchronized via ROS and stored in ROSbag files with precise timestamps, ensuring consistent multi-modal alignment for cross-sensor calibration studies.
IV-B Deep Learning Model Training
IV-B1 Data Annotation and Dataset Generation
We developed a multi-sensor annotation toolkit for efficiently labeling common objects in both camera images and point cloud data. It combines automatic and manual annotation strategies to balance speed and labeling quality. In camera images, a YOLO-based algorithm automatically generates bounding boxes, which can be manually refined. For LiDAR data, background extraction and DBSCAN clustering detect objects, producing preliminary bounding boxes that are also subject to manual adjustment. Once detections from both sensors are complete, the toolkit provides a dual-view interface to match identical objects across modalities. Using this system, Dataset 1 (1200 frames) had 800 frames annotated for a total of 5815 identical objects, while Dataset 2 (600 frames) had 200 frames annotated for 619 identical objects.
IV-B2 Training Details
@@
All deep learning models were trained from scratch on the UArizona High-Performance Computing Platform, which featured a single Nvidia 32GB V100S GPU, an AMD Zen2 processor with 5 cores, and 30 GB of RAM. Training used PyTorch 2.0 with the Adam optimizer (momentum of 0.937, weight decay of $5\times 10^{-4}$), a cosine learning rate schedule starting at 0.001 and decaying to 0.00001, and a 0.05 warm-up ratio. An exponential moving average (EMA) with a decay rate of 0.9999 was applied for added stability. Both datasets were split into training, validation, and test sets, with 90% used for training and validation and the remaining 10% reserved for testing. Within the training–validation split, 90% was allocated for training and 10% for validation, and no data augmentation was used other than resizing. The Common Feature Discriminator was trained for 300 epochs with a batch size of 4, while the Attention-based Model was trained for 800 epochs with a batch size of 8 and a token length of 256.
IV-C Results and Discussion
IV-C1 Coarse Calibration with Common Feature
@@
Common Feature Discriminator Performance
Table I summarizes the Common Feature Discriminator’s performance on both datasets, revealing consistently strong binary classification accuracies (98.00% and 92.50% for Datasets 1 and 2, respectively). These high scores indicate that the discriminator is highly effective at distinguishing whether pairs of LiDAR and camera detections correspond to the same physical object. Meanwhile, the slightly lower image and LiDAR classification accuracies reflect the inherent variability in each modality’s appearance and point cloud density, as well as the increased complexity of Dataset 2’s urban traffic scenes. Overall, the results confirm that the discriminator robustly balances object-level matching (binary classification) with semantic recognition (modality-specific class labels). Fig. 6 further illustrates the model’s qualitative behavior: two distinct objects (“differ”) are correctly identified as different, while two identical objects from different sensor views are consistently classified as “same.” This underlines the model’s robustness when handling variations in object types and poses. Notably, though occasional misclassifications occur—such as trucks being predicted as cars—these errors are relatively rare and do not significantly affect the system’s ability to produce reliable point correspondences.
TABLE II: Performance Comparison of Coarse Calibration and Other Methods
@@
Coarse Calibration Accuracy
Once the Common Feature Discriminator identifies matching objects across LiDAR and camera modalities and the corresponding point pairs are derived (as shown in Fig. 8(a)), Homography calibration is employed to obtain a coarse calibration matrix. To comprehensively evaluate the accuracy of this coarse solution, we compare it with several existing calibration methods: Manual calibration, LCCNet [1], CalibDepth [24], and CalibNet [34]. It is worth noting that for Manual calibration, we selected 34 representative point pairs uniformly distributed across the sensors’ fields of view through manual object matching. Due to the time-intensive nature of this process, we did not exhaustively select all possible point pairs. Thus, the manual calibration results presented here represent a typical calibration effort within a reasonable timeframe, rather than a full-effort exhaustive manual optimization. Table II presents the results in terms of the reprojection error metrics $\mathcal{E}_{\text{AED}}$ and $\mathcal{E}_{\text{RMSE}}$ (defined in Section III-A). From Table II, our coarse calibration demonstrates comparable or, in some cases, superior performance compared to other methods. Specifically, the automated coarse calibration outperforms Manual calibration on Dataset 2, although it exhibits a slightly higher reprojection error than Manual calibration on Dataset 1 (as visualized in Fig. 8(b)). Nevertheless, the significant advantages of the automated approach in real-time operation and reduced human intervention render this trade-off both acceptable and practical. Furthermore, while the coarse method occasionally exhibits slightly higher errors than certain deep learning–based solutions (e.g., LCCNet), it consistently surpasses others (e.g., CalibDepth and CalibNet), underscoring the effectiveness of the proposed strategy. Fig. 7 presents example calibration outcomes, with red dots (camera detections) and green dots (LiDAR detections) projected onto the image plane. Despite some minor misalignments—particularly in the upper portion of the scene (highlighted by the yellow ellipse)—the coarse calibration overall provides a notably tighter alignment between the two sensor views, potentially enabling precise LiDAR-camera fusion in real-world traffic scenarios.
TABLE III: Reprojection Error Evolution in Iterative Refinement for Dataset 1
@@
IV-C2 Fine Calibration with Iterative Refinement
@@
TABLE IV: Reprojection Error Evolution in Iterative Refinement for Dataset 2
@@
Tables III and IV detail the reprojection error evolution (using the $\mathcal{E}_{\text{AED}}$ metric) at different frame intervals (with an interval of 100 frames in our implementation) for Datasets 1 and 2. In each interval, the algorithm determines whether the newly computed homography matrix (New) provides a tighter alignment than the previously best-known matrix (Best); if so, it updates the calibration accordingly. Fig. 9(a) and 9(b) visualize these updates, where the blue line denotes the error obtained from the newly recalibrated matrix in each iteration, and the orange line tracks the evolving best-known solution. Not every recalibration step yields an improvement—reflecting the inherent noise and variability of real-world data—but key frame intervals (e.g., 0–100 for Dataset 1 and 0–1000 for Dataset 2) demonstrate significant error reductions, confirming that the iterative approach converges toward a more accurate solution over time. These updates demonstrate the iterative optimization process’s ability to adaptively refine the calibration as additional data and correspondences become available, ultimately enabling the iterative refinement to achieve significantly higher accuracy compared to the initial coarse calibration (as shown in Fig. 11(b)).
TABLE V: Performance Comparison of Iterative Refined Calibration and Other Methods
@@
Table V compares the final refined calibration performance with the aforementioned methods. Notably, the iterative refinement outperforms manual calibration by a sizeable margin in both datasets, reducing $\mathcal{E}_{\text{AED}}$ from 131.04 to 95.89 in Dataset 1 and from 40.79 to 28.35 in Dataset 2. It also consistently surpasses CalibDepth and CalibNet, while maintaining a competitive edge against LCCNet. These results demonstrate the effectiveness of iteratively incorporating new point correspondences in mitigating decalibrations and refining the sensor alignment. In practice, the procedure offers a compelling balance between accuracy, adaptability, and reduced reliance on strictly supervised or fully manual calibration protocols—making it especially valuable in long-term deployment scenarios.
Figure 14: Improved results with Attention-based Refinement over Iterative Refinement.
@@
IV-C3 Fine Calibration with Attention-based Refinement
Although the iterative refinement approach already demonstrates strong performance, it remains inherently limited by the planar assumptions of Homography. Our proposed attention-based refinement aims to mitigate errors caused by image distortions and non-planar surfaces. As shown in Table VI, calibration after applying attention-based refinement (i.e., CalibRefine) achieves lower reprojection errors than other methods on both datasets, surpassing the iterative refinement (Table V) in most metrics. Fig. 11(c) offers a more granular view of these improvements by comparing the $\mathcal{E}_{\text{AED}}$ distributions of iterative refinement and attention-based refinement. While the latter still exhibits some overlap with the former, its overall distribution skews toward smaller errors, indicating a more consistently accurate alignment between LiDAR and camera data. Fig. 14 visually illustrates such performance gains of attention-based refinement over iterative refinement.
TABLE VI: Performance Comparison of CalibRefine and Other Methods
@@ IV-C3 Fine Calibration with Attention-based Refinement
Overall, our proposed CalibRefine framework consolidates three core components—Coarse Calibration, Iterative Refinement, and Attention-Based Refinement—into a unified solution. As illustrated in Fig. 11, each stage progressively refines the LiDAR–camera alignment, mitigating errors introduced by imperfect correspondence matching (coarse stage), limited point redundancy (iterative stage), or planar homography assumptions (attention-based stage). Table VI further demonstrates that CalibRefine surpasses existing state-of-the-art methods in terms of quantitative reprojection accuracy. Beyond numerical metrics, Fig. 15 and 16 offer visual validation on Datasets 1 and 2, respectively, revealing how CalibRefine more reliably overlays LiDAR points with their corresponding image objects—particularly at scene edges and larger distances. In addition, Fig. 17 examines the distribution of calibration errors ($\mathcal{E}_{\text{AED}}$ and $\mathcal{E}_{\text{RMSE}}$) across competing approaches. Not only does CalibRefine exhibit a lower median error, but the overall spread of high-error outliers is also reduced, indicating its consistent performance. These findings underscore the robustness and adaptability of CalibRefine in real-world traffic environments.
V Conclusion
In this paper, we presented CalibRefine, an end-to-end, fully automatic, targetless, and online LiDAR–camera calibration framework that integrates three core steps—coarse calibration, iterative refinement, and attention-based refinement—into a unified pipeline. By combining robust object detection with a Common Feature Discriminator, our method circumvents the need for manually placed fiducials or human-labeled sensor parameters. The coarse calibration phase provides a strong initial alignment, which the iterative refinement then continuously improves by leveraging newly acquired point correspondences across frames. Finally, the attention-based stage applies a Vision Transformer and cross-attention to handle non-planar distortions and subtle mismatches beyond the scope of homography. Experiments on real-world urban datasets confirm that CalibRefine achieves accurate sensor alignment comparable to, and often better than, existing methods. Moving forward, the approach could benefit from exploring more advanced deep learning architectures or sophisticated mapping mechanisms, as well as extending the attention mechanism to incorporate scene geometry. Such enhancements could enable even more precise and high-fidelity calibration, particularly in large-scale deployment scenarios.